problem: the gasoline tank on a car can be a box with dimensions 1.5 ft x 2.5 ft x 9 inches. How many gallons of gasoline does it hold and what is the weight of the gasoline in a full tank?
I am trying to help my daughter with this and I do not know how to figure out formula for weight

Answer by ankor@dixie-net.com(22740)

problem: the gasoline tank on a car can be a box with dimensions 1.5 ft x 2.5 ft x 9 inches. How many gallons of gasoline does it hold and what is the weight of the gasoline in a full tank?
:
First we find the volume of the tank in cu/ft
Convert 9 in: 9/12 = .75 ft
V = 1.5 * 2.5 * .75 = 2.8125 cu/ft
:
Convert cu/ft to gallons: (1 cu/ft = 7.481 gals)
2.8125 * 7.481 = 21.04 gallons
:
One gallon of gas weighs about 6.1 lbs (lighter than water which weighs about 8.3 lb per gal)
21.04 * 6.1 = 128.346 lbs, a full tank of gas.